The cliff divers of Acapulco amaze tourists with their diving skills. The velocity v of a diver is a function of the time t the diver has fallen. If v = 15 feet per second when t = 0 seconds and v = 40 feet per second when t = 5 seconds, express v as a function of time t.
V=

Answer by stanbon(75887)

If v = 15 feet per second when t = 0 seconds and v = 40 feet per second when t = 5 seconds, express v as a function of time t.
-----
You have two points: (0,15) and (5,40)
-----
Assuming that v and t are related linearly form a linear equation.
-----
slope = (40-15)/(5-0) = 25/5 = 5
----
intercept = V(0) = 15
------
Equation:

V = 5t + 15
--------------------
------------------------
Cheers,
Stan H.